# Equilibrium Investment with Random Risk Aversion: (Non-)uniqueness, Optimality, and Comparative Statics

### v2

@@ 5 Comparative Statics
Recalling that $Q_{i}(x)=\dfrac{\delta_{i}(1-p)e^{-\delta_{i}\frac{x}{2}}}{p+(1-p)e^{-\delta_{i}\frac{x}{2}}}$, we know that $Q_{i}$ is strictly decreasing and satisfies $U_{i}=U_{i}(p)\triangleq Q_{i}(0)=\delta_{i}(1-p)$. Moreover, $\Lambda(t)=\int_{0}^{v_{i}(t)}(r_{0}+Q_{i}(x))^{2}\,\mathrm{d}x=2\int_{Q_{i}(v_{i}(t))}^{U_{i}(p)}\dfrac{(r_{0}+q)^{2}}{q(\delta_{i}-q)}\,\mathrm{d}q\triangleq 2\int_{Q_{i}(v_{i}(t))}^{U_{i}(p)}G(q,\delta)\,\mathrm{d}q$ where the second equality arises from $q=Q_{i}(x)$, $\mathrm{d}x=\dfrac{2}{q(q-\delta_{i})}\,\mathrm{d}q$ and $G(q,\delta)\triangleq\dfrac{(r_{0}+q)^{2}}{q(\delta-q)}$. At time $t^{*}$, let $Q=Q(p)\triangleq Q_{1}(v_{1}(t^{*}))=Q_{2}(v_{2}(t^{*}))$ and $\Lambda^{*}=\Lambda^{*}(p)\triangleq\Lambda(t^{*})$. The condition $|a_{1}(t^{*})|=|a_{2}(t^{*})|$ implies
$\Lambda^{*}=2\int_{Q}^{U_{1}(p)}G(q,\delta_{1})\,dq=2\int_{Q}^{U_{2}(p)}G(q,\delta_{2})\,dq.$ (5.1)
Differentiating (5.1) with respect to $p$ and we have
$\dfrac{1}{2}\dfrac{\mathrm{d}\Lambda^{*}}{\mathrm{d}p}$ $=G(U_{1},\delta_{1})\dfrac{\mathrm{d}U_{1}}{\mathrm{d}p}-G(Q,\delta_{1})\dfrac{\mathrm{d}Q}{\mathrm{d}p}$
$=-\delta_{1}G(U_{1},\delta_{1})-G(Q,\delta_{1})\dfrac{\mathrm{d}Q}{\mathrm{d}p}.$
Similarly, for the second distribution, we have
$\dfrac{1}{2}\dfrac{\mathrm{d}\Lambda^{*}}{\mathrm{d}p}=-\delta_{2}G(U_{2},\delta_{2})-G(Q,\delta_{2})\dfrac{\mathrm{d}Q}{\mathrm{d}p}.$
Eliminating $\dfrac{\mathrm{d}Q}{\mathrm{d}p}$ from the system, we obtain
$\dfrac{1}{2}\dfrac{\mathrm{d}\Lambda^{*}}{\mathrm{d}p}=\dfrac{\delta_{2}G(U_{2},\delta_{2})G(Q,\delta_{1})-\delta_{1}G(U_{1},\delta_{1})G(Q,\delta_{2})}{G(Q,\delta_{2})-G(Q,\delta_{1})}.$ (5.2)
Substituting $U_{i}=\delta_{i}(1-p)$ into $G$, we have
$\delta_{i}G(U_{i},\delta_{i})=\dfrac{(r_{0}+\delta_{i}(1-p))^{2}}{p(1-p)\delta_{i}}.$
Recalling the form of $Q_{i}$ and the fact that $Q_{i}$ is strictly decreasing, we have $\delta_{1}>\delta_{2}>U_{2}>Q$, which implies $\delta_{1}-Q>\delta_{2}-Q>0$ and $G(Q,\delta_{1})<G(Q,\delta_{2})$. Thus, the denominator of (5.2) is positive. For the numerator, we have
$\dfrac{\delta_{1}G(U_{1},\delta_{1})}{G(Q,\delta_{1})}$ $=\dfrac{(r_{0}+\delta_{1}(1-p))^{2}}{p(1-p)\delta_{1}}\cdot\dfrac{Q(\delta_{1}-Q)}{(r_{0}+Q)^{2}}=\dfrac{Q}{p(1-p)(r_{0}+Q)^{2}}(r_{0}+\delta_{1}(1-p))^{2}\left(1-\dfrac{Q}{\delta_{1}}\right),$
$\dfrac{\delta_{2}G(U_{2},\delta_{2})}{G(Q,\delta_{2})}$ $=\dfrac{Q}{p(1-p)(r_{0}+Q)^{2}}(r_{0}+\delta_{2}(1-p))^{2}\left(1-\dfrac{Q}{\delta_{2}}\right).$
Because $\delta_{1}>\delta_{2}$, it holds that $(r_{0}+\delta_{1}(1-p))^{2}>(r_{0}+\delta_{2}(1-p))^{2}$ and $1-\dfrac{Q}{\delta_{1}}>1-\dfrac{Q}{\delta_{2}}$. Therefore, $\dfrac{\delta_{1}G(U_{1},\delta_{1})}{G(Q,\delta_{1})}>\dfrac{\delta_{2}G(U_{2},\delta_{2})}{G(Q,\delta_{2})}$, which implies $\delta_{2}G(U_{2},\delta_{2})G(Q,\delta_{1})-\delta_{1}G(U_{1},\delta_{1})G(Q,\delta_{2})<0$. Consequently, $\dfrac{\mathrm{d}\Lambda^{*}}{\mathrm{d}p}<0$. Because $\Lambda(t)$ is a strictly decreasing function of $t$, a decrease in the required energy $\Lambda^{*}$ corresponds to an increase in the intersection time $t^{*}$. ∎
@@
Building on this analysis, we now relax the symmetry assumption to examine the individual effects. The following corollary presents the comparative statics of the intersection time with respect to the probability parameter of each investor.
@@
Under the conditions of Proposition 5.3 and Proposition 5.4, we have the following assertions.
  1. (1)
    If $p_{2}$ is fixed, $t^{*}$ is strictly increasing with respect $p_{1}$.
  2. (2)
    If $p_{1}$ is fixed, $t^{*}$ is strictly decreasing with respect $p_{2}$
@@
We retain the notations from the proof of Proposition 5.4.
Differentiating (5.1) with respect to $p_{1}$ yields
$\dfrac{1}{2}\dfrac{\partial\Lambda^{*}}{\partial p_{1}}$ $=-\delta_{1}G(U_{1},\delta_{1})-G(Q,\delta_{1})\dfrac{\partial Q}{\partial p_{1}},$
$\dfrac{1}{2}\dfrac{\partial\Lambda^{*}}{\partial p_{1}}$ $=-G(Q,\delta_{2})\dfrac{\partial Q}{\partial p_{1}}.$
Eliminating $\dfrac{\partial Q}{\partial p_{1}}$, we obtain
$\dfrac{1}{2}\dfrac{\partial\Lambda^{*}}{\partial p_{1}}=\dfrac{-\delta_{1}G(U_{1},\delta_{1})G(Q,\delta_{2})}{G(Q,\delta_{2})-G(Q,\delta_{1})}.$
Because the denominator is positive and the numerator is negative, we have $\dfrac{\partial\Lambda^{*}}{\partial p_{1}}<0$. Given that $\Lambda(t)$ is strictly decreasing, $t^{*}$ must increase.
Differentiating (5.1) with respect to $p_{2}$ yields
$\dfrac{1}{2}\dfrac{\partial\Lambda^{*}}{\partial p_{2}}$ $=-G(Q,\delta_{1})\dfrac{\partial Q}{\partial p_{2}},$
$\dfrac{1}{2}\dfrac{\partial\Lambda^{*}}{\partial p_{2}}$ $=-\delta_{2}G(U_{2},\delta_{2})-G(Q,\delta_{2})\dfrac{\partial Q}{\partial p_{2}}.$
Eliminating $\dfrac{\partial Q}{\partial p_{2}}$, we have
$\dfrac{1}{2}\dfrac{\partial\Lambda^{*}}{\partial p_{2}}=\dfrac{-\delta_{2}G(U_{2},\delta_{2})G(Q,\delta_{1})}{G(Q,\delta_{1})-G(Q,\delta_{2})}.$
Because the denominator is negative while the numerator remains negative, we have $\dfrac{\partial\Lambda^{*}}{\partial p_{2}}>0$, which implies that $t^{*}$ must decrease. ∎
It is important to note that the conditions in Proposition 5.3 and Proposition 5.4 are compatible with first-order stochastic dominance. One can construct risk aversion distributions such that $\bm{R}_{1}\succeq_{1}\bm{R}_{2}$ holds while the conditions in Proposition 5.3 and 5.4 are simultaneously satisfied. For instance, the parameters in Figure 1 satisfy these conditions with $r_{0}=1$, $\delta_{1}=4$, $\delta_{2}=1$, $\tilde{\mathbb{E}}[\bm{R}_{1}]=3>1.5=\tilde{\mathbb{E}}[\bm{R}_{2}]$ and $\lambda\equiv 0.4$. Consequently, first-order stochastic dominance fails to rule out the reversal of investment magnitudes characterized by $|\pi_{1}(0,T)|>|\pi_{2}(0,T)|$ for a sufficiently large $T$ and we can also observe the single crossing property.
However, we demonstrate that the desired monotonicity can be restored under a stronger stochastic order. Specifically, we introduce the concept of reverse hazard rate order and show that if $\bm{R_{1}}$ dominates $\bm{R_{2}}$ in this order, the intuition holds, i.e., $|\pi_{1}(t)|\leq|\pi_{2}(t)|$.
Definition 5.7.
Let $\bm{R_{1}}$ and $\bm{R_{2}}$ be two random variables with cumulative distribution functions $F_{\bm{R_{1}}}$ and $F_{\bm{R_{2}}}$, respectively. We say that $\bm{R_{1}}$ dominates $\bm{R_{2}}$ in the reverse hazard rate order, denoted as $\bm{R_{1}}\succeq_{rh}\bm{R_{2}}$, if the ratio
$\frac{F_{\bm{R_{2}}}(\gamma)}{F_{\bm{R_{1}}}(\gamma)}$
is decreasing in $\gamma$ for all $\gamma$ in the union of the supports of $\bm{R_{1}}$ and $\bm{R_{2}}$ where $F_{\bm{R_{1}}}(\gamma)>0$.
The following proposition compares the equilibrium strategies of two investors ranked by the reverse hazard rate order.
@@
Let $\bm{R_{1}}$ and $\bm{R_{2}}$} be two non-negative random variables with $\tilde{\mathbb{E}}\left[\bm{R_{i}}\right]<+\infty$ ($i=1,2$), such that $\bm{R_{1}}\succeq_{rh}\bm{R_{2}}$ and their distributions are distinct. Let $\pi_{i}(t)=(\sigma^{\top})^{-1}a_{i}(t)$ be the corresponding deterministic equilibrium strategies. For any $t\in[0,T)$, the magnitude of $a_{i}(t)$ satisfies
$|a_{1}(t)|\leq|a_{2}(t)|.$
Moreover, if $\lambda(t)\neq 0$, the inequality is strict. As a result, the investor with higher risk aversion adopts a less aggressive equilibrium strategy. i.e., $|\pi_{1}(t)|\leq|\pi_{2}(t)|$.
@@
Because $h_{1}(x)<h_{2}(x)$ for all positive $s$, it follows that $H_{1}(y)>H_{2}(y)$ for all $y>0$. This implies that the inverse functions satisfy $H_{2}^{-1}(z)<H_{2}^{-1}(z)$ for all $z>0$. From Theorem 3.2 (1), we have $v_{i}(t)=H_{i}^{-1}(\Lambda(t))$, where $\Lambda(t)=\int_{t}^{T}|\lambda(s)|^{2}ds$ and $v_{i}(t)=\int_{t}^{T}|a_{i}(s)|^{2}\mathrm{d}s$. Consequently, monotonicity of the inverse functions implies $v_{1}(t)\leq v_{2}(t)$. By (2.5), we have $|a_{i}(t)|=h(v_{i}(t),\bm{R_{i}})|\lambda(t)|$. From Proposition A.6, we know that $h_{1}(v_{1}(t))<h_{2}(v_{1}(t))$. Because $v_{1}(t)\leq v_{2}(t)$, it holds that $h_{2}(v_{1}(t))\leq h_{2}(v_{2}(t))$. Combining these relations yields $h_{1}(v_{1}(t))<h_{2}(v_{2}(t))$. Multiplying by $|\lambda(t)|$, we obtain $|a_{1}(t)|\leq|a_{2}(t)|$. If $\lambda(t)\neq 0$, the strict inequality $|a_{1}(t)|<|a_{2}(t)|$ is preserved. ∎
For the case where the expected risk aversion is infinite, assume that the optimal equilibria exist for both $\bm{R}_{1}$ and $\bm{R}_{2}$. Let $a_{i}(t)$ denote the deterministic coefficient corresponding to the optimal equilibrium strategy for $\bm{R}_{i}$ ($i=1,2$). Note that $\varphi(0)=\min\{t\in[0,T]\mid\Lambda(t)=0\}$ is determined solely by the market parameters and is therefore identical for both investors. Consequently, we can establish the comparative result by examining the magnitudes of $a_{i}(t)$ on the intervals $0,\varphi(0))$ and $\varphi(0),T]$ separately. Specifically, on the interval $0,\varphi(0))$, the analysis parallels the finite expectation case derived in Proposition 5.8, and thus the monotonicity of the equilibrium strategies is preserved. On the interval $\varphi(0),T]$, the definition of $\varphi(0)$ implies that $\lambda(t)=0$ almost everywhere. Consequently, the equilibrium strategies for both investors vanish, trivially satisfying the comparison inequality. Based on this decomposition, we have the following corollary.
Corollary 5.9.
Let $\bm{R}_{1}$ and $\bm{R}_{2}$ be two non-negative random variables with $\tilde{\mathbb{E}}[\bm{R}_{i}]=\infty$ ($i=1,2$), such that $\bm{R}_{1}\succeq_{rh}\bm{R}_{2}$ and their distributions are distinct. Assume that the optimal equilibria exist for both investors (i.e., $H_{i}(\infty)>\Lambda(0)$, $i=1,2$). Let $\pi_{i}(t)$ be the corresponding deterministic optimal equilibrium strategies. Then, for any $t\in[0,T)$, the magnitude of $\pi_{i}(t)$ satisfies
$|\pi_{1}(t)|\leq|\pi_{2}(t)|.$
Moreover, if $\lambda(t)\neq 0$, the inequality is strict.
While the previous results focus on comparing distinct distributions ranked by stochastic orders, it is also of theoretical interest to understand the effect of preference aggregation. We now investigate the equilibrium strategy when the risk aversion parameter is formed by a convex combination (i.e., a weighted average) of independent and identically distributed components.
Proposition 5.10.
Let $\bm{R}_{1},\cdots,\bm{R}_{n}$ be i.i.d. non-negative random variables with $\tilde{\mathbb{E}}\left[\bm{R_{i}}\right]<+\infty$ ($i=1,2,\cdots,n$) and $\bm{R}=\sum_{i=1}^{n}w_{i}\bm{R}_{i}$ be a convex combination with $w_{i}\in(0,1)$ and $\sum_{i=1}^{n}w_{i}=1$. Let $a(t)$ and $a_{i}(t)$ denote the risk exposure vector associated to the risk aversion $\bm{R}$ and $\bm{R}_{i}$, respectively. Then we have
$|a(t)|\leq|a_{1}(t)|\quad\text{for all }t\in[0,T).$
If $\bm{R}_{i}$ is non-degenerate and $\lambda(t)\neq 0$, the inequality is strict.
Proof.
Let $h(x)$ and $v(t)$ (resp. $h_{i}(x)$ and $v_{i}(t)$) be the functions associated to $\bm{R}$ (resp. $\bm{R_{i}}$ ). From the definition of $\bm{R}$, we have
$\dfrac{1}{h(x)}=\sum_{j=1}^{n}w_{j}\dfrac{\tilde{\mathbb{E}}[\bm{R}_{j}e^{-\frac{1}{2}w_{j}\bm{R}_{j}x}]}{\tilde{\mathbb{E}}[e^{-\frac{1}{2}w_{j}\bm{R}_{j}x}]}$ (5.3)
Because $\bm{R}_{j}$ are identically distributed to $\bm{R}_{1}$, each fraction in the sum is exactly $\dfrac{1}{h_{1}(w_{j}x)}$. Thus (5.3) is actually
$\dfrac{1}{h(x)}=\sum_{j=1}^{n}w_{j}\dfrac{1}{h_{1}(w_{j}x)}$
According to Proposition A.6, we know that $\dfrac{1}{h_{1}(x)}$ is decreasing in $x$. Because $w_{j}\in(0,1)$ and $x>0$, we have $w_{j}x<x$ and $\dfrac{1}{h_{1}(w_{j}x)}\geq\dfrac{1}{h_{1}(x)}$ and
$\dfrac{1}{h(x)}=\sum_{j=1}^{n}w_{j}\dfrac{1}{h_{1}(w_{j}x)}\geq\sum_{j=1}^{n}w_{j}\dfrac{1}{h_{1}(x)}=\dfrac{1}{h_{1}(x)}$
This implies $h(x)\leq h_{1}(x)$. Analogous to the proof in Proposition 5.8, we have $v(t)\leq v_{1}(t)$. Finally, because $h(x)$ is increasing in $x$, we have
$|a(t)|=h(v(t))|\lambda(t)|\leq h(v_{1}(t))|\lambda(t)|\leq h_{1}(v_{1}(t))|\lambda(t)|=|a_{1}(t)|$
If $\bm{R}_{i}$ is non-degenerate and $\lambda(t)\neq 0$, $h$ is strictly increasing, and all inequalities in the above proof become strict. Thus we have $|a(t)|<|a_{1}(t)|$
It is important to note that the inequality established above relies critically on the i.i.d. assumption. The following example demonstrates that if we retain independence but relax the identical distribution assumption, the equilibrium strategy for the convex combination need not be bounded by the strategies of the individual components.
Example 5.11.
Assume that there is one stock whose market price of risk is constant, $\lambda(\cdot)\equiv\lambda=0.5$. The time horizon $T=50$. The risk-aversion distributions of the two investors are:
Investor 1: $\tilde{\mathbb{P}}(\bm{R}_{1}=0.1)=0.2$, $\tilde{\mathbb{P}}(\bm{R}_{1}=8)=0.8$.
Investor 2: $\tilde{\mathbb{P}}(\bm{R}_{2}=1.5)=1$.
Assume that $\bm{R}_{1}$ and $\bm{R}_{2}$ are independent. We consider a composite investor with risk aversion $\bm{R}=0.5\bm{R}_{1}+0.5\bm{R}_{2}$.
Figure 2 displays the comparative evolution of $|a(t)|$, $|a_{1}(t)|$ and $|a_{2}(t)|$. The red-shaded Reversal Region highlights that the investor with the mixed risk aversion $\bm{R}$ adopts a strictly more aggressive position than both Investor 1 and Investor 2 when the time to maturity is sufficiently long. This counterexample shows that without the i.i.d. assumption, the magnitudes of equilibrium investment strategies for a convex combination of risk aversions need not be bounded by the strategies of the individual components.
Figure 2: Comparative evolution of $a(t)$ for a convex combination of risk aversions. The black dashed line corresponds to Investor 1, with distribution $\tilde{\mathbb{P}}(\bm{R}_{1}=0.1)=0.2,\tilde{\mathbb{P}}(\bm{R}_{1}=8.0)=0.8$; the blue dash-dotted line corresponds to Investor 2, with degenerate distribution $\tilde{\mathbb{P}}(\bm{R}_{2}=1.5)=1.0$. The red solid line corresponds to the mixed investor with risk aversion $\bm{R}=0.5\bm{R}_{1}+0.5\bm{R}_{2}$. The red-shaded Reversal Region indicates where $|a(t)|>\max\{|a_{1}(t)|,|a_{2}(t)|\}$. Parameters: $\lambda=0.5$, horizon $T=50$.
We now specifically investigate the aggregation of i.i.d. risk aversions. Distinct from the general result above, the following proposition utilizes the i.i.d. property to prove that the equilibrium strategy converges monotonically to its limit.
Proposition 5.12.
Suppose that $\{\bm{R}_{i}\}_{i=1}^{\infty}$ is a sequence of independent and identically distributed non-negative random variables with $\tilde{\mathbb{E}}\left[\bm{R_{i}}\right]=\mu\in(0,+\infty)$ ($i=1,2,\cdots$,). Let $S_{n}=\frac{1}{n}\sum_{i=1}^{n}\bm{R}_{i}$ denote the sample mean risk aversion and $a_{n}(t)$ denote the risk exposure vector associated to the risk aversion $S_{n}$. Then, as $n\to\infty$, we have
$|a_{n}(t)|\searrow\frac{|\lambda(t)|}{\mu}\quad\text{for all }t\in[0,T).$ (5.4)
Proof.
Let $h_{n}(x)$ and $v_{n}(t)$ be the functions associated to $S_{n}$. First, substituting $S_{n}=\frac{1}{n}\sum_{i=1}^{n}\bm{R}_{i}$ into the definition of $h_{n}(x)$ and utilizing the i.i.d. property of $\{\bm{R}_{i}\}$, we obtain
$h_{n}(x)=\frac{\tilde{\mathbb{E}}\left[\prod_{i=1}^{n}e^{-\frac{x}{2n}\bm{R}_{i}}\right]}{\displaystyle\frac{1}{n}\sum_{j=1}^{n}\tilde{\mathbb{E}}\left[\bm{R}_{j}e^{-\frac{x}{2n}\bm{R}_{j}}\prod_{k\neq j}e^{-\frac{x}{2n}\bm{R}_{k}}\right]}$ $=\frac{(\tilde{\mathbb{E}}\left[e^{-\frac{x}{2n}\bm{R}_{1}}\right])^{n}}{\tilde{\mathbb{E}}\left[\bm{R}_{1}e^{-\frac{x}{2n}\bm{R}_{1}}\right](\tilde{\mathbb{E}}\left[e^{-\frac{x}{2n}\bm{R}_{1}}\right])^{n-1}}$
$=\frac{\tilde{\mathbb{E}}\left[e^{-\frac{1}{2}\bm{R}_{1}(\frac{x}{n})}\right]}{\tilde{\mathbb{E}}\left[\bm{R}_{1}e^{-\frac{1}{2}\bm{R}_{1}(\frac{x}{n})}\right]}=h_{1}\left(\frac{x}{n}\right).$
According to Proposition A.6, $h_{1}(x)$ is increasing. For any fixed $x\geq 0$, the sequence $\frac{x}{n}$ is decreasing in $n$, implying that $h_{n}(x)$ is decreasing in $n$. Furthermore, because $H_{n}(y)=\int_{0}^{y}h_{n}^{-2}(z)dz$ and $v_{n}(t)=H_{n}^{-1}(\Lambda(t))$, the monotonicity of $h_{n}$ implies that $v_{n}(t)$ is strictly decreasing in $n$. Then we have
$h_{n+1}(v_{n+1}(t))\leq h_{n}(v_{n+1}(t))\leq h_{n}(v_{n}(t)),$
equivalently, we have
$|a_{n+1}(t)|\leq|a_{n}(t)|,\quad\forall t\in[0,T).$
Now we determine the limit of $|a_{n}(t)|$. By the Strong Law of Large Numbers, $S_{n}\to\mu$ almost surely as $n\to\infty$. This implies that $S_{n}$ converges in distribution to the degenerate random variable $R\equiv\mu$. Moreover, because $\tilde{\mathbb{E}}[S_{n}]=\mu$ for all $n$, the assumptions of Proposition 3.8 are satisfied. Thus we conclude that
$\lim_{n\to\infty}|a_{n}(t)|=h(v(t))|\lambda(t)|=\frac{|\lambda(t)|}{\mu}.$
Combining this with the monotonicity established above, we obtain (5.4). ∎

### v3

@@ 5.1 Two-Point Distributed RRA and Single-Crossing
See Appendix B.5. ∎
The condition $\dfrac{\mathrm{d}t^{*}}{\mathrm{d}p}>0$ indicates that the crossing time to maturity is strictly decreasing with respect to the probability parameter $p$. This confirms that a higher common probability of the smallest risk aversion value brings the crossing closer to maturity.
@@
Building on the analysis in Proposition 5.6, we now relax the assumption $p_{1}=p_{2}$ and investigate the individual effects of the probability parameters $p_{i}$ ($i=1,2$). The following proposition establishes the comparative statics of the crossing time w.r.t. each investor’s probability parameter.
Proposition 5.7.
Let $T>0$ be fixed. Assume that the risk aversions $\bm{R}_{1}$ and $\bm{R}_{2}$ satisfy (5.1) with $\tilde{\mathbb{E}}[\bm{R}_{1}]>\tilde{\mathbb{E}}[\bm{R}_{2}]$ and that $\lambda(t)\neq 0$ for all $t\in[0,T)$ with $\lambda(T-)\neq 0$. Let $a_{i,p_{i}}$ denote the risk exposure vector corresponding to $\bm{R}_{i}$ with $\tilde{\mathbb{P}}(\bm{R}_{i}=r_{0})=p_{i}$ for $i=1,2$. Let $t^{*}(p_{1},p_{2})\in(0,T)$ denote the crossing time, where $|a_{1,p_{1}}(t^{*}(p_{1},p_{2}))|=|a_{2,p_{2}}(t^{*}(p_{1},p_{2}))|$. Let
$O\triangleq\left\{(p_{1},p_{2})\in(0,1)^{2}:(1-p_{1})\delta_{1}>(1-p_{2})\delta_{2}\text{ and }t^{*}(p_{1},p_{2})\in(0,T)\right\}.$
Then, the set $O$ is open [^7] The set $O$ is nonempty provided that $\Lambda(0)$ is sufficiently large. The argument is analogous to that in Proposition 5.6. and the function $t^{*}(\cdot,\cdot)$ is continuously differentiable in $O$. Moreover, for any $(p_{1},p_{2})\in O$, the partial derivatives satisfy
$\frac{\partial t^{*}}{\partial p_{1}}(p_{1},p_{2})>0\quad\text{and}\quad\frac{\partial t^{*}}{\partial p_{2}}(p_{1},p_{2})<0.$
Proof.
See Appendix B.6. ∎
Similar to Proposition 5.6, the differentiability of $t^{*}$ established in Proposition 5.7 implies that the crossing time varies smoothly with respect to the probability parameters. Specifically, $t^{*}$ is strictly increasing in $p_{1}$ and strictly decreasing in $p_{2}$. This indicates that a higher probability of the smallest risk aversion for the first investor brings the crossing closer to maturity, whereas for the second investor, it pushes the crossing farther away from maturity.
5.2 Reverse Hazard Rate Order
While a higher RRA does not guarantee a less risky investment under first-order stochastic dominance, we demonstrate that it does guarantee such a reduction under the stronger reverse hazard rate order.
We first recall the definition of the reverse hazard rate order (see Capéraà (1988) and Shaked and Shanthikumar (2007)).
Definition 5.8.
Let $\bm{R}_{1}$} and $\bm{R}_{2}$} be two random variables with cumulative distribution functions $F_{{\bm{R}_{1}}}$ and $F_{{\bm{R}_{2}}}$, respectively. We say that $\bm{R}_{1}$} dominates $\bm{R}_{2}$} in the reverse hazard rate order, denoted ${\bm{R}_{1}}\succeq_{rh}{\bm{R}_{2}}$, if the ratio $F_{{\bm{R}_{2}}}/F_{{\bm{R}_{1}}}$ is decreasing [^8] Herein, “increasing” means “non-decreasing” and “decreasing” means “non-increasing.” in $\{\gamma:F_{{\bm{R}_{1}}}(\gamma)>0\}\cup\{\gamma:F_{{\bm{R}_{2}}}(\gamma)>0\}$. Here, we use the convention $\frac{p}{0}=\infty$ for $p>0$.
The following proposition establishes a comparison of the risk exposure magnitudes under the reverse hazard rate order.
Proposition 5.9.
Let $\bm{R}_{1}$} and $\bm{R}_{2}$} be two non-negative random variables with $\tilde{\mathbb{E}}[{\bm{R}_{i}}]<\infty$ ($i=1,2$) and $F_{\bm{R}_{1}}\neq F_{\bm{R}_{2}}$. If ${\bm{R}_{1}}\succeq_{rh}{\bm{R}_{2}}$ and $H_{2}(\infty)>\Lambda(0)$, then $|a_{1}(t)|\leq|a_{2}(t)|$ for any $t\in[0,T)$, where the strict inequality holds if $\lambda(t)\neq 0$.
Proof.
See Appendix B.7. ∎
We now consider the case where the expectation of risk aversion is infinite. Assume that the optimal equilibria exist for both $\bm{R}_{1}$ and $\bm{R}_{2}$. Note that $\varphi(0)=\min\{t\in[0,T]:\,\Lambda(t)=0\}$ is determined solely by the market parameters and is therefore identical for both investors. Consequently, the comparative result can be established by examining the magnitudes of $a_{i}(t)$ on the intervals $0,\varphi(0))$ and $\varphi(0),T]$ separately. Specifically, on the interval $0,\varphi(0))$, the analysis parallels the finite expectation case derived in Proposition 5.9, and thus the order of the risk exposure magnitude is preserved. On the interval $\varphi(0),T)$, the definition of $\varphi(0)$ implies that $\lambda\equiv\textbf{0}$. As a result, the corresponding risk exposure magnitude also satisfies the desired order. Therefore, we obtain the following corollary.
@@ 5.2 Reverse Hazard Rate Order
Let $\bm{R}_{1}$ and $\bm{R}_{2}$ be two non-negative random variables with $\tilde{\mathbb{E}}[\bm{R}_{i}]=\infty$ ($i=1,2$) and $F_{\bm{R}_{1}}\neq F_{\bm{R}_{2}}$. Assume that $\bm{R}_{1}\succeq_{rh}\bm{R}_{2}$ and the optimal equilibria exist for both investors (i.e., $H_{1}(\infty)\geq H_{2}(\infty)>\Lambda(0)$). Let $a_{i}(\cdot)$ be the risk exposure vector corresponding to the optimal equilibrium strategies. Then $|a_{1}(t)|\leq|a_{2}(t)|$ for any $t\in[0,T)$, where the strict inequality holds if $\lambda(t)\neq 0$.
5.3 Aggregation of Risk Aversions
The preceding results focus on comparisons between investments for distinct risk-aversion distributions ordered by stochastic dominance relations. We next study the effect of preference aggregation on the risk exposure magnitude. The following proposition shows that a convex combination of i.i.d. risk aversions leads to a reduction in the risk exposure magnitude.
Proposition 5.11.
Let $\bm{R}_{1},\cdots,\bm{R}_{n}$ be i.i.d. non-negative random variables with $\tilde{\mathbb{E}}\left[{\bm{R}_{1}}\right]<\infty$ and $\bm{R}\triangleq\sum_{i=1}^{n}w_{i}\bm{R}_{i}$ be a convex combination with $w_{i}\in(0,1)$ ($1\leq i\leq n$) and $\sum_{i=1}^{n}w_{i}=1$. Let $H_{1}=H_{\bm{R}_{1}}$ and $H=H_{\bm{R}}$. Assume that $H_{1}(\infty)>\Lambda(0)$. [^9] Under this assumption, the equilibrium strategy associated with each $\bm{R}_{i}$ exists and is unique by Theorem 3.2(1). Moreover, the proof in Appendix B.8 shows that $H(\infty)>\Lambda(0)$ also holds. Let $a(\cdot)$ and $a_{1}(\cdot)$ denote the risk exposure vector associated with the risk aversions $\bm{R}$ and $\bm{R}_{1}$, respectively. Then $|a(t)|\leq|a_{1}(t)|$ for all $t\in[0,T)$, where the strict inequality holds if $\bm{R}_{1}$ is non-degenerate and $\lambda(t)\neq 0$.
@@ 5.3 Aggregation of Risk Aversions
See Appendix B.8. ∎
It is important to note that the inequality established above relies critically on the i.i.d. assumption. The following example demonstrates that, if independence is retained but the identical distribution assumption is relaxed, the conclusion of Proposition 5.11 may no longer hold.
Example 5.12.
Assume that there is only one stock, whose market price of risk is constant, $\lambda(\cdot)\equiv\lambda=0.5$. The time horizon $T=50$. The risk-aversion distributions of the two investors are:
  • Investor 1: $\tilde{\mathbb{P}}(\bm{R}_{1}=0.1)=0.2$ and $\tilde{\mathbb{P}}(\bm{R}_{1}=8)=0.8$;
  • Investor 2: $\tilde{\mathbb{P}}(\bm{R}_{2}=1.5)=1$.
Obviously, $\bm{R}_{1}$ and $\bm{R}_{2}$ are independent. Consider an aggregated investor whose risk aversion is given by $\bm{R}=0.5\bm{R}_{1}+0.5\bm{R}_{2}$. Figure 2 illustrates the comparative evolution of the risk exposure magnitudes $|a(\cdot)|$, $|a_{1}(\cdot)|$, and $|a_{2}(\cdot)|$. The red-shaded Reversal Region highlights that the investor with the aggregated risk aversion $\bm{R}$ adopts a strictly more aggressive position than both Investor 1 and Investor 2 when the time to maturity is sufficiently long.
Figure 2: Comparative evolution of risk exposure magnitudes for individual risk aversions and their convex combination. The black dashed line corresponds to Investor 1 with $\tilde{\mathbb{P}}(\bm{R}_{1}=0.1)=0.2$ and $\tilde{\mathbb{P}}(\bm{R}_{1}=8)=0.8$; the blue dash-dotted line corresponds to Investor 2 with degenerate distribution $\tilde{\mathbb{P}}(\bm{R}_{2}=1.5)=1$; the red solid line corresponds to the aggregated investor with risk aversion $\bm{R}=0.5\bm{R}_{1}+0.5\bm{R}_{2}$. The red-shaded Reversal Region indicates where $|a(t)|>\max\{|a_{1}(t)|,|a_{2}(t)|\}$. Parameters: $\lambda=0.5$, $T=50$.
Proposition 5.11 establishes a comparison result for convex combinations of i.i.d. risk aversions with a fixed population size, but does not address the asymptotic behavior as the number of samples increases. We therefore focus on the special case of the sample mean risk aversion. The following proposition examines the limit of the risk exposure vector as the sample size $n$ tends to infinity and establishes monotonic convergence by exploiting the structure of the sample mean.
@@
Suppose that $\{\bm{R}_{i}\}_{i=1}^{\infty}$ is a sequence of i.i.d. non-negative random variables with $\tilde{\mathbb{E}}\left[{\bm{R}_{1}}\right]=\mu\in(0,\infty)$. Let $\bm{Y}_{n}\triangleq\frac{1}{n}\sum_{i=1}^{n}\bm{R}_{i}$ denote the sample mean risk aversion. Let $H_{1}$ (resp. $\tilde{H}_{n}$) denote the function defined in (3.2) corresponding to $\bm{R}_{1}$(resp. $\bm{Y}_{n}$). Assume that $H_{1}(\infty)>\Lambda(0)$. Then the deterministic equilibrium strategies associated with each $\bm{R}_{i}$} and $\bm{Y}_{n}$ exist and are unique. Let $\tilde{a}_{n}(\cdot)$ be the risk exposure vector associated to the risk aversion $\bm{Y}_{n}$. Then, as $n\to\infty$, we have
$|\tilde{a}_{n}(t)|\searrow\frac{|\lambda(t)|}{\mu}\quad\text{for all }t\in[0,T).$ (5.2)
@@
See Appendix B.9. ∎